OpenStudy (anonymous):

6xsquared - y=98, find x if y = -2

OpenStudy (theviper):

\[6x^2-y=98\]\[6x^2-(-2)=98\]\[6x^2+2=98\]\[6x^2=98-2=96\]\[x^2=96/6=16\]\[x=4\]

OpenStudy (anonymous):

Substitute in y \[6x^2 - (-2) = 98 \\ 6x^2 = 96 \\ x^2 = 16 \\ x = \pm 4 \]

OpenStudy (lgbasallote):

^looks like someone's wrong.....wonder which one

Parth (parthkohli):

@TheViper p/m 4.

OpenStudy (lgbasallote):

^sure?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Don't forget that if you root something, there's always a possibility of a negative answer! :P My grade 9 teacher used to say we should factor: \[x^2-16 = 0 \\ (x+4)(x-4) = 0 \\ x=4,-4\] But my awesome grade 10/11/12 honors teacher lets us get away with it. :)

Parth (parthkohli):

@muntoo is the right guy. :D

OpenStudy (lgbasallote):

no one's the engaging guy though.........

Parth (parthkohli):

Agreed

OpenStudy (lgbasallote):

i wonder if it's the adrenaline or users nowadays just love to race...

OpenStudy (theviper):

@muntoo is right so go with @muntoo :)

OpenStudy (anonymous):

Over at StackExchange, we call this the "Fastest Gun in the West" syndrome (FGITW). The theory is that "fast guns" give quick little hints so they deserve credit as much as the slow and thorough answerers.

OpenStudy (anonymous):

awesome +4-4 is right!!! TY

OpenStudy (lgbasallote):

strange coincidence...our name for it here has something to do with guns too.. "answer sniping"

OpenStudy (anonymous):

+4 nd - 4

OpenStudy (anonymous):

+-4